### v3

@@ Abstract
In the past several years, the last-iterate convergence of the Stochastic Gradient Descent (SGD) algorithm has triggered people’s interest due to its good performance in practice but lack of theoretical understanding. For Lipschitz convex functions, different works have established the optimal $O(\log(1/\delta)\log T/\sqrt{T})$ or $O(\sqrt{\log(1/\delta)/T})$ high-probability convergence rates for the final iterate, where $T$ is the time horizon and $\delta$ is the failure probability. However, to prove these bounds, all the existing works are either limited to compact domains or require almost surely bounded noise. It is natural to ask whether the last iterate of SGD can still guarantee the optimal convergence rate but without these two restrictive assumptions. Besides this important question, there are still lots of theoretical problems lacking an answer. For example, compared with the last-iterate convergence of SGD for non-smooth problems, only few results for smooth optimization have yet been developed. Additionally, the existing results are all limited to a non-composite objective and the standard Euclidean norm. It still remains unclear whether the last-iterate convergence can be provably extended to wider composite optimization and non-Euclidean norms. In this work, to address the issues mentioned above, we revisit the last-iterate convergence of stochastic gradient methods and provide the first unified way to prove the convergence rates both in expectation and in high probability to accommodate general domains, composite objectives, non-Euclidean norms, Lipschitz conditions, smoothness, and (strong) convexity simultaneously. Additionally, we extend our analysis to obtain the last-iterate convergence under heavy-tailed noise.
@@ 1 Introduction
In this paper, we consider the constrained composite optimization problem $\min_{x\in\mathcal{X}}F(x)\coloneqq f(x)+h(x)$ where both $f(x)$ and $h(x)$ are convex (but possibly satisfying additional conditions such as strong convexity, smoothness, etc.) and $\mathcal{X}\subseteq\mathbb{R}^{d}$ is a nonempty closed convex set. Since a true gradient is computationally prohibitive to obtain (e.g., large-scale machine learning tasks) or even infeasible to access (e.g., streaming data), the classic Stochastic Gradient Descent (SGD) (Robbins and Monro, 1951) algorithm has emerged to be the gold standard for a light-weight yet effective computational procedure commonly adopted in production for the majority of machine learning tasks: SGD only requires a stochastic first-order oracle $\widehat{\partial}f(x)$ satisfying $\mathbb{\mathbb{E}}[\widehat{\partial}f(x)\mid x]\in\partial f(x)$ where $\partial f(x)$ denotes the set of subgradients at $x$ and guarantees provable convergence under certain conditions (e.g., Lipschitz condition for $f(x)$ and finite variance on the stochastic oracle).
@@ 1.1 Our Contributions
  • We prove the last iterate of CSMD can converge in the rate of $O(1/\sqrt{T})$ for smooth convex optimization and $\widetilde{O}(1/T)$ for smooth strongly convex problems both in expectation and in high probability for any general domain $\mathcal{X}$, hence resolving Q2.
@@ 2 Preliminaries
Notations: $\mathbb{\mathbb{N}}$ is the set of natural numbers (excluding 0). $\left[d\right]\coloneqq\left\{1,2,\cdots,d\right\}$ for any $d\in\mathbb{\mathbb{N}}$. $a\lor b$ and $a\land b$ are defined as $\max\left\{a,b\right\}$ and $\min\left\{a,b\right\}$, respectively. $\left\langle\cdot,\cdot\right\rangle$ is the standard Euclidean inner product on $\mathbb{R}^{d}$. $\left\|\cdot\right\|$ represents a general norm on $\mathbb{R}^{d}$ and $\left\|\cdot\right\|_{*}$ is its dual norm. Given a set $A\subseteq\mathbb{R}^{d}$, $\mathrm{int}(A)$ stands for its interior points. For a function $f$, $\partial f(x)$ denotes the set of subgradients at $x$.
@@
    Unbiased gradient estimator: For a given $x^{t}\in\mathcal{X}$ in the $t$-th iteration, we can access an unbiased gradient estimator $\widehat{g}^{t}$, i.e., $\mathbb{\mathbb{E}}\left[\widehat{g}^{t}\mid\mathcal{F}^{t-1}\right]\in\partial f(x^{t})$, where $\mathcal{F}^{t}\coloneqq\sigma\left(\widehat{g}^{s},s\in\left[t\right]\right)$ is the natural filtration.
@@
    Finite variance: $\exists\sigma\geq 0$ denoting the noise level such that $\mathbb{\mathbb{E}}\left[\left\|\xi^{t}\right\|_{*}^{2}\mid\mathcal{F}^{t-1}\right]\leq\sigma^{2}$ where $\xi^{t}\coloneqq\widehat{g}^{t}-\mathbb{\mathbb{E}}\left[\widehat{g}^{t}\mid\mathcal{F}^{t-1}\right]$.
@@
    Sub-Gaussian noise: $\exists\sigma\geq 0$ denoting the noise level such that $\mathbb{\mathbb{E}}\left[\exp\left(\lambda\left\|\xi^{t}\right\|_{*}^{2}\right)\mid\mathcal{F}^{t-1}\right]\leq\exp\left(\lambda\sigma^{2}\right),\forall\lambda\in\left[0,\sigma^{-2}\right]$.
@@
Given a sigma algebra $\mathcal{F}$ and a random vector $Z\in\mathbb{R}^{d}$ that is $\mathcal{F}$-measurable, if $\xi\in\mathbb{R}^{d}$ is a random vector satisfying $\mathbb{\mathbb{E}}\left[\xi\mid\mathcal{F}\right]=0$ and $\mathbb{\mathbb{E}}\left[\exp\left(\lambda\left\|\xi\right\|_{*}^{2}\right)\mid\mathcal{F}\right]\leq\exp\left(\lambda\sigma^{2}\right),\forall\lambda\in\left[0,\sigma^{-2}\right]$, then
$\mathbb{\mathbb{E}}\left[\exp\left(\left\langle\xi,Z\right\rangle\right)\mid\mathcal{F}\right]\leq\exp\left(\sigma^{2}\left\|Z\right\|^{2}\right).$
@@ 3 Last-Iterate Convergence of Stochastic Gradient Methods
The algorithm, Composite Stochastic Mirror Descent, is presented in Algorithm 1. When $h(x)=0$, Algorithm 1 degenerates to the standard Stochastic Mirror Descent algorithm. If we further consider the case $\left\|\cdot\right\|=\left\|\cdot\right\|_{2}$, Algorithm 1 can recover the standard projected SGD by taking $\psi(x)=\frac{1}{2}\left\|x\right\|_{2}^{2}$. We assume $T\geq 2$ throughout the following paper to avoid some algebraic issues in the proof. The full version of every following theorem with its proof is deferred into the appendix.
@@ 3.1 General Convex Functions
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x^{*})\right]\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{T}+\frac{(M+\sigma)\sqrt{D_{\psi}(x^{*},x^{1})}\log T}{\sqrt{T}}\right).$
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x^{*})\right]\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{T}+\frac{(M+\sigma)\sqrt{D_{\psi}(x^{*},x^{1})\log T}}{\sqrt{T}}\right).$
@@
In Theorem 3.3, we present the high-probability bounds for $(L,M)$-smooth functions. Again, the constant $\eta$ is picked to get the best dependence on the parameters $M,\sigma,D_{\psi}(x^{*},x^{1})$ and $\log(1/\delta)$. The full version of Theorem 3.3 with arbitrary $\eta$, Theorem C.2, is deferred into the appendix. Compared with Theorem 3.1, the high-probability rates only incur an extra $O(\sqrt{\log(1/\delta)})$ factor (or $O(\log(1/\delta))$ for arbitrary $\eta$, which is known to be optimal for $L=0$ (Harvey et al., 2019b)).
@@
3.1.1 Optimal Rates via the Linear Decay Step Size
In this section, we show that the $O(\sqrt{\log T})$ factor appeared in Theorems 3.1 and 3.3 for the case of known $T$ can be totally removed. Even restricted to the special case of Lipschitz objectives (i.e., $L=0$ in Assumption 3.), this is a strict improvement compared to Jain et al. (2021), in which the $O(1/\sqrt{T})$ rate is obtained under the bounded noise assumption on a compact domain $\mathcal{X}$.
@@ 3.1.1 Optimal Rates via the Linear Decay Step Size
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x^{*})\right]\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{T}+\frac{(M+\sigma)\sqrt{D_{\psi}(x^{*},x^{1})}}{\sqrt{T}}\right).$
@@
The step size decaying linearly was introduced by Zamani and Glineur (2025), in which the authors proved that it ensures the last iterate of the subgradient method converges at the rate $O(1/\sqrt{T})$ for non-composite non-smooth convex objectives with the Euclidean norm. Here, we show it can also guarantee the optimal $O(L/T+(M+\sigma)/\sqrt{T})$ rate in our setting. Again, the constant $\eta$ can be set to any positive number. The full statements of Theorems 3.4 and 3.5 are provided in Appendix C.1.
@@ 3.2 Strongly Convex Functions
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x^{*})\right]\leq\begin{cases}O\left(\frac{LD_{\psi}(x^{*},x^{1})}{T}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{f}(T+\kappa_{f})}\right)&\eta_{t\in\left[T\right]}=\frac{1}{\mu_{f}(t+2\kappa_{f})}\\ O\left(\frac{L(1+\kappa_{f})D_{\psi}(x^{*},x^{1})}{T(T+\kappa_{f})}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{f}(T+\kappa_{f})}\right)&\eta_{t\in\left[T\right]}=\frac{2}{\mu_{f}(t+1+4\kappa_{f})}\end{cases}.$
If $T$ is known, by taking $\eta_{t\in\left[T\right]}=\begin{cases}\frac{1}{\mu_{f}(1+2\kappa_{f})}&t=1\\ \frac{1}{\mu_{f}(\eta+2\kappa_{f})}&2\leq t\leq\tau\\ \frac{2}{\mu_{f}(t-\tau+2+4\kappa_{f})}&t\geq\tau+1\end{cases}$ with $\eta\coloneqq 1.5$ and $\tau\coloneqq\left\lceil\frac{T}{2}\right\rceil$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x^{*})\right]\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{\exp\left(\frac{T}{3+4\kappa_{f}}\right)}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{f}(T+\kappa_{f})}\right).$
The in-expectation rates are stated in Theorem 3.6, where the constant $\eta=1.5$ is chosen without any special reason. Generally speaking, it can be any non-negative number that satisfies $\eta+\kappa_{f}>1$. The interested reader could refer to Theorem D.1 in the appendix for a completed version of Theorem 3.6. We would like to remind the reader that $\kappa_{f}\geq 1$ is not necessary, as we are considering the general $(L,M)$-smooth functions. Hence, it can be zero.
As before, we first take $L=0$ to consider the special Lipschitz case. Due to $\kappa_{f}=0$ now, all bounds will degenerate to $O(\log T/T)$, which is known to be optimal for the step size $\nicefrac{{1}}{{\mu_{f}t}}$ (Harvey et al., 2019b) and only incurs an extra $O(\log T)$ factor compared with the best $O(1/T)$ bound when $T$ is known (Jain et al., 2021). We would also like to mention that Theorem 3.6 is the first to give the in-expectation last-iterate bound for the step size $\nicefrac{{2}}{{\mu_{f}(t+1)}}$. Interestingly, the extra $O(\log T)$ factor appears again compared to the known $O(1/T)$ bound on the function value gap for the non-uniform averaging strategy under this step size (Lacoste-Julien et al., 2012). Besides, Lacoste-Julien et al. (2012) also shows $\mathbb{\mathbb{E}}\left[\left\|x^{T+1}-x^{*}\right\|_{2}^{2}\right]=O(1/T)$. Whereas it is currently unknown whether our $\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x^{*})\right]=O(\log T/T)$ bound can be improved to match the $O(1/T)$ rate.
@@
To finish this section, we provide the high-probability convergence results in Theorem 3.7. Again, the constant $\eta=1.5$ is set without any particular reason. The full statement with general $\eta$, Theorem D.2, can be found in the appendix. Besides, $\kappa_{f}$ is possible to be zero as mentioned above. Compared with Theorem 3.6, only an additional $O(\log(1/\delta))$ factor appears. Such an extra loss is known to be inevitable for $L=0$ due to Harvey et al. (2019b).
For the Lipschitz case (i.e., $L=\kappa_{f}=0$), by noticing $D_{\psi}(x^{*},x^{1})=O(M^{2}/\mu_{f}^{2})$ [^7] This holds now due to $\mu_{f}\left\|x^{*}-x^{1}\right\|^{2}/2\leq\mu_{f}D_{\psi}(x^{*},x^{1})\leq M\left\|x^{*}-x^{1}\right\|$., all of these bounds will degenerate to $O(\log(1/\delta)\log T/T)$ matching the best-known last-iterate bound proved by Harvey et al. (2019b) for the step size $\nicefrac{{1}}{{\mu_{f}t}}$. For the step size $\nicefrac{{2}}{{\mu_{f}(t+1)}}$, Harvey et al. (2019a) has proved the high-probability bound $O(\log(1/\delta)/T)$ for the non-uniform averaging output instead of the last iterate. Hence, as far as we know, our high-probability rate for the step size $\nicefrac{{2}}{{\mu_{f}(t+1)}}$ is new. However, we would like to mention that our bound for known $T$ is worse by a logarithmic factor than that of Jain et al. (2021), though, which assumes bounded noise.

### v4

@@ Abstract
In the past several years, the last-iterate convergence of the Stochastic Gradient Descent (SGD) algorithm has triggered people’s interest due to its good performance in practice but lack of theoretical understanding. For Lipschitz convex functions, different works have established the optimal $O(\log(1/\delta)\log T/\sqrt{T})$ or $O(\sqrt{\log(1/\delta)/T})$ high-probability convergence rates for the final iterate, where $T$ is the time horizon and $\delta$ is the failure probability. However, to prove these bounds, all the existing works are either limited to compact domains or require almost surely bounded noise. It is natural to ask whether the last iterate of SGD can still guarantee the optimal convergence rate but without these two restrictive assumptions. Besides this important question, there are still lots of theoretical problems lacking an answer. For example, compared with the last-iterate convergence of SGD for non-smooth problems, only few results for smooth optimization have yet been developed. Additionally, the existing results are all limited to a non-composite objective and the standard Euclidean norm. It still remains unclear whether the last-iterate convergence can be provably extended to wider composite optimization and non-Euclidean norms. In this work, to address the issues mentioned above, we revisit the last-iterate convergence of stochastic gradient methods and provide the first unified way to prove the convergence rates both in expectation and in high probability to accommodate general domains, composite objectives, non-Euclidean norms, Lipschitz conditions, smoothness, and (strong) convexity simultaneously. Additionally, we extend our analysis to obtain last-iterate convergence under heavy-tailed and sub-Weibull noise.
@@ 1 Introduction
In this paper, we consider the constrained composite optimization problem, $\min_{x\in\mathcal{X}}F(x)\coloneqq f(x)+h(x)$, where both $f(x)$ and $h(x)$ are convex (but possibly satisfying additional conditions such as strong convexity, smoothness, etc.) and $\mathcal{X}\subseteq\mathbb{R}^{d}$ is a nonempty closed convex set. Since a true gradient is computationally prohibitive to obtain (e.g., large-scale machine learning tasks) or even infeasible to access (e.g., streaming data), the classic Stochastic Gradient Descent (SGD) (Robbins and Monro, 1951) algorithm has emerged to be the gold standard for a light-weight yet effective computational procedure commonly adopted in production for the majority of machine learning tasks: SGD only requires a stochastic first-order oracle $\widehat{\partial}f(x)$ satisfying $\mathbb{E}[\widehat{\partial}f(x)\mid x]\in\partial f(x)$ where $\partial f(x)$ denotes the set of subgradients at $x$ and guarantees provable convergence under certain conditions (e.g., Lipschitz condition for $f(x)$ and finite variance on the stochastic oracle).
@@ 1.1 Our Contributions
  • We prove the last iterate of CSMD can converge in the rate of $O(1/\sqrt{T})$ for smooth convex optimization and $O(1/T)$ for smooth strongly convex problems both in expectation and in high probability for any general domain $\mathcal{X}$, hence resolving Q2.
@@ 2 Preliminaries
Notations: $\mathbb{N}$ is the set of natural numbers (excluding 0). $\left[d\right]\coloneqq\left\{1,2,\cdots,d\right\}$ for any $d\in\mathbb{N}$. $a\lor b$ and $a\land b$ are defined as $\max\left\{a,b\right\}$ and $\min\left\{a,b\right\}$, respectively. $\left\langle\cdot,\cdot\right\rangle$ is the standard Euclidean inner product on $\mathbb{R}^{d}$. $\left\|\cdot\right\|$ represents a general norm on $\mathbb{R}^{d}$ and $\left\|\cdot\right\|_{*}$ is its dual norm. Given a set $A\subseteq\mathbb{R}^{d}$, $\mathrm{int}(A)$ stands for its interior points. For a function $f$, $\partial f(x)$ denotes the set of subgradients at $x$.
@@
    Unbiased gradient estimator: For a given $x^{t}\in\mathcal{X}$ in the $t$-th iteration, we can access an unbiased gradient estimator $\widehat{g}^{t}$, i.e., $\mathbb{E}\left[\widehat{g}^{t}\mid\mathcal{F}^{t-1}\right]\in\partial f(x^{t})$, where $\mathcal{F}^{t}\coloneqq\sigma\left(\widehat{g}^{s},s\in\left[t\right]\right)$ is the natural filtration.
@@
    Finite variance: $\exists\sigma\geq 0$ denoting the noise level such that $\mathbb{E}\left[\left\|\xi^{t}\right\|_{*}^{2}\right]\leq\sigma^{2}$ where $\xi^{t}\coloneqq\widehat{g}^{t}-\mathbb{E}\left[\widehat{g}^{t}\mid\mathcal{F}^{t-1}\right]$.
@@
    Sub-Gaussian noise: $\exists\sigma\geq 0$ denoting the noise level such that $\mathbb{E}\left[\exp\left(\lambda\left\|\xi^{t}\right\|_{*}^{2}\right)\mid\mathcal{F}^{t-1}\right]\leq\exp\left(\lambda\sigma^{2}\right),\forall\lambda\in\left[0,\sigma^{-2}\right]$.
@@
Given a sigma algebra $\mathcal{F}$ and a random vector $Z\in\mathbb{R}^{d}$ that is $\mathcal{F}$-measurable, if $\xi\in\mathbb{R}^{d}$ is a random vector satisfying $\mathbb{E}\left[\xi\mid\mathcal{F}\right]=0$ and $\mathbb{E}\left[\exp\left(\lambda\left\|\xi\right\|_{*}^{2}\right)\mid\mathcal{F}\right]\leq\exp\left(\lambda\sigma^{2}\right),\forall\lambda\in\left[0,\sigma^{-2}\right]$, then
$\mathbb{E}\left[\exp\left(\left\langle\xi,Z\right\rangle\right)\mid\mathcal{F}\right]\leq\exp\left(\sigma^{2}\left\|Z\right\|^{2}\right).$
@@ 3 Last-Iterate Convergence of Stochastic Gradient Methods
The algorithm, Composite Stochastic Mirror Descent, is presented in Algorithm 1. When $h(x)=0$, Algorithm 1 degenerates to the standard Stochastic Mirror Descent algorithm. If we further consider the case $\left\|\cdot\right\|=\left\|\cdot\right\|_{2}$, Algorithm 1 can recover the standard projected SGD by taking $\psi(x)=\frac{1}{2}\left\|x\right\|_{2}^{2}$. We assume $T\geq 2$ throughout the following paper to avoid some algebraic issues in the proof. The full version of every following theorem with its proof is deferred to the appendix.
@@ 3.1 General Convex Functions
$\mathbb{E}\left[F(x^{T+1})-F(x^{*})\right]\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{T}+\frac{(M+\sigma)\sqrt{D_{\psi}(x^{*},x^{1})}\log T}{\sqrt{T}}\right).$
@@
$\mathbb{E}\left[F(x^{T+1})-F(x^{*})\right]\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{T}+\frac{(M+\sigma)\sqrt{D_{\psi}(x^{*},x^{1})\log T}}{\sqrt{T}}\right).$
@@
In Theorem 3.3, we present the high-probability bounds for $(L,M)$-smooth functions. Again, the constant $\eta$ is picked to get the best dependence on the parameters $M,\sigma,D_{\psi}(x^{*},x^{1})$ and $\log(1/\delta)$. The full version of Theorem 3.3 with arbitrary $\eta$, Theorem C.2, is deferred to the appendix. Compared with Theorem 3.1, the high-probability rates only incur an extra $O(\sqrt{\log(1/\delta)})$ factor (or $O(\log(1/\delta))$ for arbitrary $\eta$, which is known to be optimal for $L=0$ (Harvey et al., 2019b)).
@@
3.1.1 Optimal Rates via the Linearly Decaying Step Size
In this section, we show that the $O(\sqrt{\log T})$ factor appearing in Theorems 3.1 and 3.3 for known $T$ can be completely removed. Even restricted to the special case of Lipschitz objectives (i.e., $L=0$ in Assumption 3.), this is a strict improvement compared to Jain et al. (2021), in which the $O(1/\sqrt{T})$ rate is obtained under the bounded noise assumption on a compact domain $\mathcal{X}$.
@@ 3.1.1 Optimal Rates via the Linearly Decaying Step Size
$\mathbb{E}\left[F(x^{T+1})-F(x^{*})\right]\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{T}+\frac{(M+\sigma)\sqrt{D_{\psi}(x^{*},x^{1})}}{\sqrt{T}}\right).$
@@
The linearly decaying step size was introduced by Zamani and Glineur (2025), in which the authors proved that it ensures the last iterate of the subgradient method converges at the rate $O(1/\sqrt{T})$ for non-composite non-smooth convex objectives with the Euclidean norm. Here, we show it can also guarantee the optimal $O(L/T+(M+\sigma)/\sqrt{T})$ rate in our setting. Again, the constant $\eta$ can be set to any positive number. The full statements of Theorems 3.4 and 3.5 are provided in Appendix C.1.
@@ 3.2 Strongly Convex Functions
$\mathbb{E}\left[F(x^{T+1})-F(x^{*})\right]\leq\begin{cases}O\left(\frac{LD_{\psi}(x^{*},x^{1})}{T}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{f}(T+\kappa_{f})}\right)&\eta_{t\in\left[T\right]}=\frac{1}{\mu_{f}(t+2\kappa_{f})}\\ O\left(\frac{L(1+\kappa_{f})D_{\psi}(x^{*},x^{1})}{T(T+\kappa_{f})}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{f}(T+\kappa_{f})}\right)&\eta_{t\in\left[T\right]}=\frac{2}{\mu_{f}(t+1+4\kappa_{f})}\end{cases}.$
If $T$ is known, by taking $\eta_{t\in\left[T\right]}=\begin{cases}\frac{1}{\mu_{f}(1+2\kappa_{f})}&t\leq\tau\\ \frac{2}{\mu_{f}(t-\tau+2+4\kappa_{f})}&t\geq\tau+1\end{cases}$ with $\tau\coloneqq\left\lceil\frac{T}{2}\right\rceil$, there is
$\mathbb{E}\left[F(x^{T+1})-F(x^{*})\right]\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{\exp\left(\frac{T}{2+4\kappa_{f}}\right)}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{f}(T+\kappa_{f})}\right).$
The in-expectation rates are stated in Theorem 3.6. We would like to remind the reader that $\kappa_{f}\geq 1$ is not necessary, as we are considering the general $(L,M)$-smooth functions. Hence, it can be zero.
As before, we first take $L=0$ to consider the special Lipschitz case. Due to $\kappa_{f}=0$ now, all bounds will degenerate to $O(\log T/T)$, which is known to be optimal for the step size $\nicefrac{{1}}{{\mu_{f}t}}$ (Harvey et al., 2019b) and only incurs an extra $O(\log T)$ factor compared with the best $O(1/T)$ bound when $T$ is known (Jain et al., 2021). We would also like to mention that Theorem 3.6 is the first to give the in-expectation last-iterate bound for the step size $\nicefrac{{2}}{{\mu_{f}(t+1)}}$. Interestingly, the extra $O(\log T)$ factor appears again compared to the known $O(1/T)$ bound on the function value gap for the non-uniform averaging strategy under this step size (Lacoste-Julien et al., 2012). Besides, Lacoste-Julien et al. (2012) also shows $\mathbb{E}\left[\left\|x^{T+1}-x^{*}\right\|_{2}^{2}\right]=O(1/T)$. Whereas it is currently unknown whether our $\mathbb{E}\left[F(x^{T+1})-F(x^{*})\right]=O(\log T/T)$ bound can be improved to match the $O(1/T)$ rate under the step size $\nicefrac{{2}}{{\mu_{f}(t+1)}}$.
@@
Now, we provide the high-probability convergence results in Theorem 3.7. The constant $\eta=1.5$ is set without any particular reason. The full statement with general $\eta$, Theorem D.2, can be found in the appendix. Besides, $\kappa_{f}$ is possible to be zero as mentioned above. Compared with Theorem 3.6, only an additional $O(\log(1/\delta))$ factor appears. Such an extra loss is known to be inevitable for $L=0$ due to Harvey et al. (2019b).
For the Lipschitz case (i.e., $L=\kappa_{f}=0$), by noticing $D_{\psi}(x^{*},x^{1})=O(M^{2}/\mu_{f}^{2})$ [^7] This holds now due to $\mu_{f}\left\|x^{*}-x^{1}\right\|^{2}/2\leq\mu_{f}D_{\psi}(x^{*},x^{1})\leq M\left\|x^{*}-x^{1}\right\|$., all of these bounds will degenerate to $O(\log(1/\delta)\log T/T)$ matching the best-known last-iterate bound proved by Harvey et al. (2019b) for the step size $\nicefrac{{1}}{{\mu_{f}t}}$. For the step size $\nicefrac{{2}}{{\mu_{f}(t+1)}}$, Harvey et al. (2019a) has proved the high-probability bound $O(\log(1/\delta)/T)$ for the non-uniform averaging output instead of the last iterate. Hence, as far as we know, our high-probability rate for the step size $\nicefrac{{2}}{{\mu_{f}(t+1)}}$ is new.
@@
3.2.1 Optimal Rates via a New Step Size
In this section, we show that the $O(\log T)$ factor appearing in Theorems 3.6 and 3.7 for known $T$ can be completely removed. Even restricted to the special case of Lipschitz objectives (i.e., $L=0$ in Assumption 3.), this is a strict improvement compared to Jain et al. (2021), in which the $O(1/T)$ rate is obtained under the bounded noise assumption on a compact domain $\mathcal{X}$ (though $\mathcal{X}$ has to be bounded now, our proof does not rely on this fact). Moreover, to avoid some algebraic issues in the proof, we assume $T\geq 4$ in this part.
Theorem 3.8.
Under Assumptions 1.-4. and 5A. with $\mu_{f}>0$ and $\mu_{h}=0$, let $\kappa_{f}\coloneqq\frac{L}{\mu_{f}}\geq 0$, if $T$ is known, by taking $\eta_{t\in\left[T\right]}=\begin{cases}\frac{1}{\mu_{f}(1+2\kappa_{f})}&t\leq\tau_{1}\\ \frac{2}{\mu_{f}(t-\tau_{1}+2+4\kappa_{f})}&\tau_{1}+1\leq t\leq\tau_{2}\\ \frac{T-t+1}{2L(T-\tau_{2})}\land\frac{T-t+1}{\mu_{f}(T-\tau_{2})(T+\kappa_{f})}&t\geq\tau_{2}+1\end{cases}$ with $\tau_{1}\coloneqq\left\lceil\frac{T}{4}\right\rceil$ and $\tau_{2}\coloneqq\left\lceil\frac{T}{2}\right\rceil$, there is
$\mathbb{E}\left[F(x^{T+1})-F(x^{*})\right]\leq O\left(\frac{LD_{\psi}(x^{*},x^{1})}{\exp\left(\frac{T}{4+8\kappa_{f}}\right)}+\frac{M^{2}+\sigma^{2}}{\mu_{f}(T+\kappa_{f})}\right).$
Theorem 3.9.
Under Assumptions 1.-4. and 5B. with $\mu_{f}>0$ and $\mu_{h}=0$, let $\kappa_{f}\coloneqq\frac{L}{\mu_{f}}\geq 0$ and $\delta\in(0,1)$, if $T$ is known, by taking $\eta_{t\in\left[T\right]}=\begin{cases}\frac{1}{\mu_{f}(\eta+2\kappa_{f})}&t\leq\tau_{1}\\ \frac{2}{\mu_{f}(t-\tau_{1}+2+4\kappa_{f})}&\tau_{1}+1\leq t\leq\tau_{2}\\ \frac{T-t+1}{2L(T-\tau_{2})}\land\frac{T-t+1}{\mu_{f}(T-\tau_{2})(T+2\kappa_{f})}&t\geq\tau_{2}+1\end{cases}$ with $\eta\coloneqq 1.5$, $\tau_{1}\coloneqq\left\lceil\frac{T}{4}\right\rceil$ and $\tau_{2}\coloneqq\left\lceil\frac{T}{2}\right\rceil$, then with probability at least $1-\delta$, there is
$F(x^{T+1})-F(x^{*})\leq O\left(\frac{\mu_{f}(1+\kappa_{f})D_{\psi}(x^{*},x^{1})}{\exp\left(\frac{T}{6+8\kappa_{f}}\right)}+\frac{M^{2}+\sigma^{2}\log\frac{1}{\delta}}{\mu_{f}(T+\kappa_{f})}\right).$
As far as we know, Theorems 3.8 and 3.9 are the first results to obtain the optimal $O(1/T)$ rate for the last iterate without any restrictive assumptions. We highlight that the step size developed here differs significantly from the schedule introduced by Jain et al. (2021). Concretely, our new step size is a combination of the classical one (as used in Theorems 3.6 and 3.7) and the linearly decaying step size proposed in Zamani and Glineur (2025). Moreover, we note that the constant $\eta=1.5$ is picked without any particular reason. The full statements of Theorems 3.8 and 3.9 are provided in Appendix D.1.1.